# InfMasking: Unleashing Synergistic Information by Contrastive Multimodal Interactions

### v2

@@ Appendix F Pseudo-Code
17:  $\mathcal{L}_{\text{InfMasking}}\leftarrow-\frac{1}{M^{\prime}}\sum\limits_{k=1}^{M^{\prime}}\Big[\mathbb{E}_{\begin{subarray}{c}{z^{\prime}}_{mask}^{k},{z^{\prime}}_{\text{pos}}\sim p({Z^{\prime}}_{mask}^{k},Z^{\prime})\end{subarray}}\left[\log\frac{\exp({{z^{\prime}}_{mask}^{k}}^{T}z^{\prime}_{\text{pos}}/\tau)}{\exp({{z^{\prime}}_{mask}^{k}}^{T}z^{\prime}_{\text{pos}}/\tau)+\sum_{z^{\prime}_{\text{neg}}}\exp({{z^{\prime}}_{mask}^{k}}^{T}z^{\prime}_{\text{neg}}/\tau)}\right]$
18:         $\mathbb{E}_{\begin{subarray}{c}{z^{\prime\prime}}_{mask}^{k},z^{\prime\prime}_{\text{pos}}\sim p({Z^{\prime\prime}}_{mask}^{k},Z^{\prime\prime})\end{subarray}}\left[\log\frac{\exp({{z^{\prime\prime}}_{mask}^{k}}^{T}z^{\prime\prime}_{\text{pos}}/\tau)}{\exp({{z^{\prime\prime}}_{mask}^{k}}^{T}z^{\prime\prime}_{\text{pos}}/\tau)+\sum_{z^{\prime\prime}_{\text{neg}}}\exp({{z^{\prime\prime}}_{mask}^{k}}^{T}z^{\prime\prime}_{\text{neg}}/\tau)}\right]\Big]$
@@ Appendix G Proof
$=\mathbb{E}_{\text{mask}}[\mathbb{E}_{z^{\prime}\sim p(Z^{\prime})}\left[\log\frac{\exp(z^{\prime T}z^{\prime}_{{\text{mask}}}/\tau)}{\exp(z^{\prime T}z^{\prime}_{{\text{mask}}}/\tau)+\sum_{z^{\prime}_{\text{neg}}}\exp(z_{neg}^{{}^{\prime}T}z^{\prime}_{{\text{mask}}}/\tau)}\right]]$ (9)
$=\mathbb{E}_{z^{\prime}\sim p(Z^{\prime})}[\mathbb{E}_{\text{mask}}\left[\log\frac{\exp(z^{\prime T}z^{\prime}_{{\text{mask}}}/\tau)}{\exp(z^{\prime T}z^{\prime}_{{\text{mask}}}/\tau)+\sum_{z^{\prime}_{\text{neg}}}\exp(z_{neg}^{{}^{\prime}T}z^{\prime}_{{\text{mask}}}/\tau)}\right]]$ (10)
$=\mathbb{E}_{z^{\prime}\sim p(Z^{\prime})}[\mathbb{E}_{\text{mask}}\left[(z^{\prime T}z^{\prime}_{{\text{mask}}}/\tau)-\log[\exp(z^{\prime T}z^{\prime}_{{\text{mask}}}/\tau)+\sum_{z^{\prime}_{\text{neg}}}\exp(z_{neg}^{{}^{\prime}T}z^{\prime}_{{\text{mask}}}/\tau)]\right]]$ (11)
$\geq\mathbb{E}_{z^{\prime}\sim p(Z^{\prime})}\left[z^{\prime~T}\mathbb{E}_{\text{mask}}[z^{\prime}_{{\text{mask}}}]/\tau-\log\mathbb{E}_{\text{mask}}[\exp(z^{\prime T}z^{\prime}_{{\text{mask}}}/\tau)+\sum_{z^{\prime}_{\text{neg}}}\exp(z_{neg}^{{}^{\prime}T}z^{\prime}_{{\text{mask}}}/\tau)]\right]$ (12)
@@
$\geq\mathbb{E}_{z^{\prime}\sim p(Z^{\prime})}\left[z^{\prime T}\boldsymbol{\mu}_{z_{\text{mask}}^{\prime}}/\tau-\log[\exp(z^{\prime T}\boldsymbol{\mu}_{z_{\text{mask}}^{\prime}}/\tau+\frac{1}{2\tau^{2}}z^{\prime T}\boldsymbol{\Sigma}_{z_{\text{mask}}^{\prime}}z)+\sum_{z^{\prime}_{\text{neg}}}\exp(z_{neg}^{{}^{\prime}T}z^{\prime}_{{\text{mask}}}/\tau)]\right]$ (15)

### v3

@@ Appendix F Pseudo-Code
17:  $\mathcal{L}_{\text{InfMasking}}\leftarrow-\frac{1}{M^{\prime}}\sum\limits_{k=1}^{M^{\prime}}\Big[\mathbb{E}_{\begin{subarray}{c}{z^{\prime}}_{mask}^{k},{z^{\prime}}_{\text{pos}}\sim p({Z^{\prime}}_{mask}^{k},Z^{\prime})\end{subarray}}\left[\log\frac{\exp({{z^{\prime}}_{mask}^{k}}^{T}z^{\prime}_{\text{pos}}/\tau)}{\exp({{z^{\prime}}_{mask}^{k}}^{T}z^{\prime}_{\text{pos}}/\tau)+\sum_{z^{\prime}_{\text{neg}}}\exp(z^{\prime T}z^{\prime}_{neg}/\tau)}\right]$
18:         $\mathbb{E}_{\begin{subarray}{c}{z^{\prime\prime}}_{mask}^{k},z^{\prime\prime}_{\text{pos}}\sim p({Z^{\prime\prime}}_{mask}^{k},Z^{\prime\prime})\end{subarray}}\left[\log\frac{\exp({{z^{\prime\prime}}_{mask}^{k}}^{T}z^{\prime\prime}_{\text{pos}}/\tau)}{\exp({{z^{\prime\prime}}_{mask}^{k}}^{T}z^{\prime\prime}_{\text{pos}}/\tau)+\sum_{z^{\prime\prime}_{\text{neg}}}\exp(z^{\prime\prime T}z^{\prime\prime}_{neg}/\tau)}\right]\Big]$
@@ Appendix G Proof
$=\mathbb{E}_{\text{mask}}[\mathbb{E}_{z^{\prime}\sim p(Z^{\prime})}\left[\log\frac{\exp(z^{\prime T}z^{\prime}_{{\text{mask}}}/\tau)}{\exp(z^{\prime T}z^{\prime}_{{\text{mask}}}/\tau)+\sum_{z^{\prime}_{\text{neg}}}\exp(z^{\prime T}z^{\prime}_{neg}/\tau)}\right]]$ (9)
$=\mathbb{E}_{z^{\prime}\sim p(Z^{\prime})}[\mathbb{E}_{\text{mask}}\left[\log\frac{\exp(z^{\prime T}z^{\prime}_{{\text{mask}}}/\tau)}{\exp(z^{\prime T}z^{\prime}_{{\text{mask}}}/\tau)+\sum_{z^{\prime}_{\text{neg}}}\exp(z^{\prime T}z^{\prime}_{neg}/\tau)}\right]]$ (10)
$=\mathbb{E}_{z^{\prime}\sim p(Z^{\prime})}[\mathbb{E}_{\text{mask}}\left[(z^{\prime T}z^{\prime}_{{\text{mask}}}/\tau)-\log[\exp(z^{\prime T}z^{\prime}_{{\text{mask}}}/\tau)+\sum_{z^{\prime}_{\text{neg}}}\exp(z^{\prime T}z^{\prime}_{neg}/\tau)]\right]]$ (11)
$\geq\mathbb{E}_{z^{\prime}\sim p(Z^{\prime})}\left[z^{\prime T}\mathbb{E}_{\text{mask}}[z^{\prime}_{{\text{mask}}}]/\tau-\log\mathbb{E}_{\text{mask}}[\exp(z^{\prime T}z^{\prime}_{{\text{mask}}}/\tau)+\sum_{z^{\prime}_{\text{neg}}}\exp(z^{\prime T}z^{\prime}_{neg}/\tau)]\right]$ (12)
@@
$\geq\mathbb{E}_{z^{\prime}\sim p(Z^{\prime})}\left[z^{\prime T}\boldsymbol{\mu}_{z_{\text{mask}}^{\prime}}/\tau-\log[\exp(z^{\prime T}\boldsymbol{\mu}_{z_{\text{mask}}^{\prime}}/\tau+\frac{1}{2\tau^{2}}z^{\prime T}\boldsymbol{\Sigma}_{z_{\text{mask}}^{\prime}}z)+\sum_{z^{\prime}_{\text{neg}}}\exp(z^{\prime T}z^{\prime}_{neg}/\tau)]\right]$ (15)